# Long-Range Pairing in the Kitaev Model: Krylov Subspace Signatures

### v2

@@ IV.2 Krylov subspace and edge sensitivity
To quantify the relative weight carried by the odd and even subsequences, we define the Krylov staggering parameter
$\boxed{\eta_{n}\equiv\ln\!\left(\frac{b_{2n-1}}{b_{2n}}\right)\quad(n\geq 2)},$ (35)
which measures the logarithmic ratio between consecutive odd and even Lanczos coefficients. The analysis window is restricted to the numerically stable portion of the recursion; in particular, we exclude all coefficients once $b_{n}\lesssim 10^{-7}$ or any other stability criterion is violated (see Appendix B.4 for the complete set of criteria). The lower bound $n\geq 2$ excludes the first recursion step, where initialization effects can dominate. When $\eta_{n}\approx 0$, the odd and even subsequences are approximately equal; deviations from zero indicate asymmetry between the two invariant subspaces. The operational definition of sign changes and the choice of analysis window are detailed in Sec. V.3.
@@
The diagnostic power of this pattern is clearest when the edge mode is well separated from the bulk continuum (e.g., when hopping and pairing strengths are comparable): the seed couples to a distinct low-energy feature at small $n$, yielding a sharp, numerically stable signature in $\eta_{n}$. When the edge and bulk gaps become comparable (e.g., when pairing dominates overwhelmingly over hopping), the qualitative correspondence between $\eta_{n}$ and the gap-based classification persists, but quantitative matching becomes numerically challenging, as the odd and even sub-chain responses nearly coincide within finite resolution. Nonetheless, $\eta_{n}$ remains qualitatively sensitive to boundary physics (see the discussion in Sec. V.3 and Appendix D for analysis of alternative seeds); the underlying BdG eigenmode structure of these regimes is analyzed in Sec. V.1 and Appendix C.
@@
In practice, $N_{\mathrm{cross}}$ depends strongly on the seed location: boundary seeds bias the recursion toward one chain end, while bulk seeds couple comparably to both ends and can produce multiple sign flips at parametrically separated depths. The mapping between $\eta_{n}$ and edge/bulk gap classification is therefore quantitatively most reliable when using a boundary seed; data for alternative seed are provided in Appendix D.
@@ V.1 Spectrum and the role of long-range pairing
Spectra at $\epsilon=1$ and $\epsilon=10$ (Appendix C) confirm these trends persist across pairing regimes, without altering the fundamental $\alpha$-driven eigenmode reorganization. Therefore, we focus henceforth on $\epsilon=-0.2$, where moderate density near $E=0$ facilitates numerical resolution of the odd- and even-parity Krylov subspaces while preserving the essential long-range physics.
@@ V.2 Lanczos Coefficients
We now present representative Lanczos-coefficient data that underlie the crossing-count diagnostic and connection to edge-versus-bulk gap physics. For the Hermitian boundary seed $\gamma_{1}$, all diagonal Lanczos coefficients vanish, $a_{n}=0$ (Appendix B.2), so the Krylov subspace representation is purely off-diagonal and is fully characterized by the sequence $\{b_{n}\}$. In this setting it is natural to view the recursion as two interleaved “sub-chains” (odd and even steps), whose relative ordering is quantified by the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ ($n\geq 2$), as introduced and discussed in Sec. IV.2. All data presented satisfy rigorous numerical stability criteria; in particular, we exclude all Lanczos coefficients once $b_{n}\lesssim 10^{-7}$ or any other stability check fails (see Appendix B.4 for details).
@@
Unless stated otherwise, the phase diagrams are computed for an open chain with $N=1000$ and a boundary seed operator, and are evaluated on a uniform $99\times 99$ grid in $(\alpha,\theta)$ where $\alpha\in(0,3]$ and $\theta\in(0,\pi)$. Boundary-localized seeds provide the sharpest quantitative agreement with the BdG-derived edge-bulk boundary (up to grid resolution and finite-size effects), whereas bulk seeds couple comparably to both edges and typically yield less sharp matching. Data for additional seeds are provided in Appendix D.
@@
The quality of this match depends strongly on the choice of seed operator. Boundary seeds, such as $\gamma_{1}$ used here, maximize the overlap with boundary-localized BdG modes and therefore provide the most sensitive and quantitatively reliable diagnostic of whether the lowest excitation scale is edge dominated or bulk dominated. Seeds that remain localized near the boundary but involve a sum of adjacent Majorana operators, e.g. $\gamma_{1}+\gamma_{2}$, retain clear qualitative sensitivity to the edge-bulk distinction, yet the quantitative agreement with the gap boundary is typically less sharp. In contrast, seeds placed deep in the bulk, such as $\gamma_{N}$ or $\gamma_{N}+\gamma_{N+1}$ (recalling that an $N$-site chain has $2N$ Majorana modes), couple comparably to left- and right-edge physics and predominantly probe bulk-extended excitations; as a result, the crossing-count signature becomes less tightly locked to the edge-bulk gap boundary and the quantitative agreement degrades further, even though qualitative trends can still be discerned. Taken together, these comparisons show that boundary seeds are essential for obtaining a robust and quantitative Krylov subspace diagnostic of boundary-versus-bulk control of the gap, while non-boundary seeds provide at best a weaker, more qualitative probe. Data for the aforementioned additional seeds are provided in Appendix D.
Figure 3: Phase diagram for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries ($N=1000$) and boundary seed $\gamma_{1}$, generating $2000$ Lanczos coefficients. As discussed in Sec. IV.2, the black region indicates parameters where the Krylov staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ exhibits nonzero robust sign changes ($N_{\mathrm{cross}}\geq 1$), while the white region corresponds to $N_{\mathrm{cross}}=0$. As discussed in Sec. IV.1, solid curves show the edge-bulk gap boundary $\Delta_{\mathrm{edge}}=\Delta_{\mathrm{bulk}}$ extracted from the BdG spectrum using the edge-weight criterion for three choices of threshold, $\omega_{\mathrm{edge}}=0.05,0.1,0.5$ (with $\ell_{\mathrm{edge}}=\sqrt{N}$). The Krylov-based and gap-based boundaries coincide within the numerical resolution of the $99\times 99$ grid in $(\alpha,\theta)$ with $\alpha\in(0,3]$ and $\theta\in(0,\pi)$. The phase boundary is robust across different $\omega_{\mathrm{edge}}$ values; discrepancies arise from grid resolution and finite-size effects, the latter being most pronounced in the strong long-range regime ($\alpha\lesssim 1$) at larger $\theta$ (see main text for further discussion).
@@ VI Conclusion and Outlook
This work establishes that Krylov subspace diagnostics extracted from operator dynamics can resolve physically distinct phases in the model for topological superconductivity, namely the long-range Kitaev chain. Specifically, we have demonstrated that Lanczos coefficients generated from local boundary operators provide a quantitative diagnostic of whether the lowest excitation gap in the long-range Kitaev chain is controlled by boundary-localized or bulk-extended modes. This finding challenges the prevailing expectation that Lanczos data in quadratic fermionic systems are trivial or featureless, and it shifts attention toward using Krylov diagnostics as probes of how low-energy excitations are localized along the chain and how the excitation gap is controlled by edge versus bulk degrees of freedom. The key technical advance enabling this result is an exact single-particle formulation of the operator Lanczos algorithm that achieves machine precision for chains containing hundreds of sites where the model considered is particularly well-conditioned for large recursion depth (also see Appendix B.4).
The central observable developed in this work is a sequence of odd–even Krylov staggering parameters, defined as the logarithmic ratio between consecutive odd and even Lanczos coefficients. We introduce and study the sign structure of this sequence, which correlates robustly with the edge-versus-bulk character of the lowest positive energy excitation across the full long-range phase diagram. This correlation arises from a bipartite Krylov structure induced by the interplay of three physical ingredients: U(1)-breaking pairing, power-law coupling decay, and open boundary conditions. When the lowest positive BdG mode is classified as edge localized according to the edge-weight criterion in Eq. (30), a boundary seed couples more strongly to that mode, and the Krylov recursion typically shows robust sign changes of $\eta_{n}$. Conversely, when the smallest positive BdG mode is bulk extended, the odd and even subsequences tend to keep a fixed ordering over the stable recursion window and no robust sign crossings occur. Importantly, this distinction persists across the short-range to long-range crossover, indicating that the crossing-count diagnostic captures robust edge-versus-bulk physics rather than being tied to fine-tuned model features.

### v3

@@ IV.2 Krylov subspace and edge sensitivity
To quantify the imbalance between the odd and even Krylov subsequences, we use the Krylov staggering parameter $\eta_{n}$ defined in Eq. (15), evaluated only within the numerically stable window and with $n\geq 2$ to suppress initialization effects. In the analytically solvable short-range balanced limit ($\epsilon=0$, $\alpha\to\infty$), $\eta_{n}$ is exactly constant and its sign is an exact topological order parameter (Sec. II.6); the long-range problem studied here asks how that theorem deforms once long-range couplings ($\alpha<\infty$) and pairing imbalance ($\epsilon\neq 0$) render $\eta_{n}$ recursion-depth dependent. The central empirical finding established in Sec. V, motivated directly by the exact short-range result, is that while the flatness of $\eta_{n}$ is lost, the sign structure of $\eta_{n}$ across $n$ continues to distinguish regimes governed predominantly by edge-localized low-energy physics from those governed by bulk-extended excitations.
The analysis window is restricted to the numerically stable portion of the recursion; in particular, we exclude all coefficients once $b_{n}\lesssim 10^{-7}$ or any other stability criterion is violated (see Appendix B.4 for the complete set of criteria). The lower bound $n\geq 2$ excludes the first recursion step, where initialization effects can dominate. When $\eta_{n}\approx 0$, the odd and even subsequences are approximately equal; deviations from zero indicate asymmetry between the two invariant subspaces. The operational definition of sign changes and the choice of analysis window are detailed in Sec. V.3.
@@
The diagnostic power of this pattern is clearest when the edge mode is well separated from the bulk continuum (e.g., when hopping and pairing strengths are comparable): the seed couples to a distinct low-energy feature at small $n$, yielding a sharp, numerically stable signature in $\eta_{n}$. When the edge and bulk gaps become comparable (e.g., when pairing dominates overwhelmingly over hopping), the qualitative correspondence between $\eta_{n}$ and the gap-based classification persists, but quantitative matching becomes numerically challenging, as the odd and even sub-chain responses nearly coincide within finite resolution. Nonetheless, $\eta_{n}$ remains qualitatively sensitive to boundary physics (see the discussion in Sec. V.3 and Appendix E for analysis of alternative seeds); the underlying BdG eigenmode structure of these regimes is analyzed in Sec. V.1 and Appendix D.
@@
In practice, $N_{\mathrm{cross}}$ depends strongly on the seed location: boundary seeds bias the recursion toward one chain end, while bulk seeds couple comparably to both ends and can produce multiple sign flips at parametrically separated depths. The mapping between $\eta_{n}$ and edge/bulk gap classification is therefore quantitatively most reliable when using a boundary seed; data for alternative seed are provided in Appendix E.
@@ V.1 Spectrum and the role of long-range pairing
Spectra at $\epsilon=1$ and $\epsilon=10$ (Appendix D) confirm these trends persist across pairing regimes, without altering the fundamental $\alpha$-driven eigenmode reorganization. Therefore, we focus henceforth on $\epsilon=-0.2$, where moderate density near $E=0$ facilitates numerical resolution of the odd- and even-parity Krylov subspaces while preserving the essential long-range physics.
@@ V.2 Lanczos Coefficients
Before turning to the long-range data, we recall the exact short-range anchor established in Sec. II.6: for the balanced short-range chain ($\alpha\to\infty$, $\epsilon=0$), the Lanczos coefficients are exactly flat in alternating pairs, $b_{2n-1}=|\mu|$ and $b_{2n}=2|t|$ for every $n$ and every system size $N$ (Eq. (14)). Equivalently, the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ is exactly constant, and its sign is an exact topological order parameter of the short-range Kitaev chain (Eq. (17)). The long-range data below show how this flat structure deforms once power-law couplings are introduced, while preserving the central physical lesson of the short-range theorem: the sign structure of $\eta_{n}$, and hence the crossing count $N_{\mathrm{cross}}$ defined in Eq. (39), continues to track whether the lowest excitation is edge-dominated or bulk-dominated across the full $(\theta,\alpha)$ phase diagram.
We now present representative Lanczos-coefficient data that underlie the crossing-count diagnostic and connection to edge-versus-bulk gap physics. For the Hermitian boundary seed $\gamma_{1}$, all diagonal Lanczos coefficients vanish, $a_{n}=0$ (Appendix B.2), so the Krylov subspace representation is purely off-diagonal and is fully characterized by the sequence $\{b_{n}\}$. In this setting it is natural to view the recursion as two interleaved “sub-chains” (odd and even steps), whose relative ordering is quantified by the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ ($n\geq 2$), as introduced in Sec. II.6 and further discussed in Sec. IV.2. All data presented satisfy rigorous numerical stability criteria; in particular, we exclude all Lanczos coefficients once $b_{n}\lesssim 10^{-7}$ or any other stability check fails (see Appendix B.4 for details).
@@
Figure 3: Phase diagram for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries ($N=1000$) and boundary seed $\gamma_{1}$, generating $2000$ Lanczos coefficients. As discussed in Sec. IV.2, the black region indicates parameters where the Krylov staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ exhibits nonzero robust sign changes ($N_{\mathrm{cross}}\geq 1$), while the white region corresponds to $N_{\mathrm{cross}}=0$. As discussed in Sec. IV.1, solid curves show the edge-bulk gap boundary $\Delta_{\mathrm{edge}}=\Delta_{\mathrm{bulk}}$ extracted from the BdG spectrum using the edge-weight criterion for three choices of threshold, $\omega_{\mathrm{edge}}=0.05,0.1,0.5$ (with $\ell_{\mathrm{edge}}=\sqrt{N}$). The central numerical result is the near coincidence of these independently constructed boundaries: the Krylov-based and gap-based boundaries coincide within the numerical resolution of the $99\times 99$ grid in $(\alpha,\theta)$ with $\alpha\in(0,3]$ and $\theta\in(0,\pi)$. The phase boundary is robust across different $\omega_{\mathrm{edge}}$ values; discrepancies arise from grid resolution and finite-size effects, the latter being most pronounced in the strong long-range regime ($\alpha\lesssim 1$) at larger $\theta$ (see main text for further discussion).
@@
Unless stated otherwise, the phase diagrams are computed for an open chain with $N=1000$ and a boundary seed operator, and are evaluated on a uniform $99\times 99$ grid in $(\alpha,\theta)$ where $\alpha\in(0,3]$ and $\theta\in(0,\pi)$. Boundary-localized seeds provide the sharpest quantitative agreement with the BdG-derived edge-bulk boundary (up to grid resolution and finite-size effects), whereas bulk seeds couple comparably to both edges and typically yield less sharp matching. Data for additional seeds are provided in Appendix E. Finite-size robustness of the extracted boundary is demonstrated in two complementary ways in Sec. V.4.
@@
The quality of this match depends strongly on the choice of seed operator. Boundary seeds, such as $\gamma_{1}$ used here, maximize the overlap with boundary-localized BdG modes and therefore provide the most sensitive and quantitatively reliable diagnostic of whether the lowest excitation scale is edge dominated or bulk dominated. Seeds that remain localized near the boundary but involve a sum of adjacent Majorana operators, e.g. $\gamma_{1}+\gamma_{2}$, retain clear qualitative sensitivity to the edge-bulk distinction, yet the quantitative agreement with the gap boundary is typically less sharp. In contrast, seeds placed deep in the bulk, such as $\gamma_{N}$ or $\gamma_{N}+\gamma_{N+1}$ (recalling that an $N$-site chain has $2N$ Majorana modes), couple comparably to left- and right-edge physics and predominantly probe bulk-extended excitations; as a result, the crossing-count signature becomes less tightly locked to the edge-bulk gap boundary and the quantitative agreement degrades further, even though qualitative trends can still be discerned. Taken together, these comparisons show that boundary seeds are essential for obtaining a robust and quantitative Krylov subspace diagnostic of boundary-versus-bulk control of the gap, while non-boundary seeds provide at best a weaker, more qualitative probe. Data for the aforementioned additional seeds are provided in Appendix E.
[…]
Figure 4: Finite-size robustness of the edge-versus-bulk BdG/Krylov diagnostics at $\epsilon=-0.2$ with boundary seed $\gamma_{1}$. 4 Joint edge-bulk gap and Krylov crossing-count phase diagram at $N=500$ on the same $99\times 99$ $(\theta,\alpha)$ grid used in Fig. 3, with $\ell_{\rm edge}=\lfloor\sqrt{N}\rfloor$; the extracted boundary is visually nearly indistinguishable from the $N=1000$ result in Fig. 3. 4 Edge weight $W_{\rm edge}$ versus $\theta/\pi$ for $\alpha=1/3,\,1,\,2$, $\varepsilon=-0.2$ and $N=500,600,700,800,900,1000$ (line style distinguishes $\alpha$ and color distinguishes NN). In the trivial regime, $W_{\rm edge}$ decreases with increasing $N$, consistent with vanishing boundary weight in the thermodynamic limit, while in the topological regime the curves collapse within visual resolution onto a common finite profile. Open circles mark parameter points with $N_{\rm cross}\geq 1$ [Eq. (39)]; their onset coincides with the onset of finite $W_{\rm edge}$, showing that the Krylov diagnostic tracks the same boundary-localized physics as the BdG criterion.
V.4 Finite-Size Scaling
To assess the stability of the results presented in Sec. V.3 with respect to system size, we perform two complementary finite-size checks at $\epsilon=-0.2$. First, we repeat the joint BdG edge-bulk gap and Krylov crossing-count phase-diagram analysis at $N=500$, keeping the boundary window $\ell_{\rm edge}=\lfloor\sqrt{N}\rfloor$ and all other operational parameters identical to those used for the $N=1000$ calculations as presented in Fig 3. Second, at fixed $\alpha=1/3,\,1,\,2$, we plot the edge weight $W_{\rm edge}$ as a function of $\theta$ for $N=500,600,700,800,900,1000$ and overlay the parameter points for which $N_{\rm cross}\geq 1$ as defined in Eq. (39).
Figure 4 summarizes these two checks. Figure 4 shows that even the worst-case comparison at $N=500$ yields a phase boundary indistinguishable from the $N=1000$ result used throughout the main text, demonstrating convergence at the system sizes studied in this work. Figure 4 shows that in the trivial regime $W_{\rm edge}$ decreases systematically with increasing $N$, consistent with vanishing boundary weight in the thermodynamic limit, whereas in the topological regime the curves for all shown system sizes collapse within visual resolution onto a common finite profile. The onset of finite $W_{\rm edge}$ coincides with the appearance of $N_{\rm cross}\geq 1$, confirming that the Krylov crossing diagnostic tracks the same boundary-localized physics as the BdG-based criterion.
@@ VI Conclusion and Outlook
This work establishes two central results. First, in the hopping-pairing-balanced short-range Kitaev chain ($\Delta=t$), the single-particle Krylov recursion seeded with the boundary Majorana $\gamma_{1}$ can be solved exactly: for every recursion depth $n$ and every system size $N$, the Lanczos coefficients satisfy $b_{2n-1}=|\mu|$, $b_{2n}=2|t|$, and $a_{n}=0$ (Appendix C). The corresponding staggering parameter $\eta_{n}\equiv\ln(b_{2n-1}/b_{2n})$ is therefore exactly constant, and its sign is an exact topological order parameter: $\eta<0$ if and only if $|\mu|<2|t|$ (topological phase with Majorana zero modes), while $\eta>0$ if and only if $|\mu|>2|t|$ (trivial phase). Thus Kitaev’s phase boundary is recovered from a purely Krylov-recursion argument, without invoking winding numbers, band topology, or bulk-gap closure.
Second, in the long-range Kitaev chain at $\epsilon=-0.2$, this exact flat short-range structure is deformed, but its physical content survives. We show that Lanczos coefficients generated from a local boundary seed provide a quantitative diagnostic of whether the lowest excitation gap is controlled by boundary-localized or bulk-extended modes across the full $(\theta,\alpha)$ phase diagram. The key observable is the sign structure of the Krylov staggering parameter $\eta_{n}$: robust sign changes correlate with edge-dominated low-energy physics, while fixed sign correlates with bulk-dominated regimes. This demonstrates that Lanczos data in quadratic fermionic systems are not merely trivial or featureless, but can encode sharp information about topology, localization, and the spatial origin of the excitation gap. Both results are enabled by an exact single-particle formulation of the operator Lanczos algorithm as derived in this work, which reduces the recursion from exponentially large operator space to a finite-dimensional linear problem and yields machine-precision data for chains of hundreds of sites.
When the lowest positive BdG mode is classified as edge localized according to the edge-weight criterion in Eq. (33), a boundary seed couples more strongly to that mode, and the Krylov recursion typically shows robust sign changes of $\eta_{n}$. Conversely, when the smallest positive BdG mode is bulk extended, the odd and even subsequences tend to keep a fixed ordering over the stable recursion window and no robust sign crossings occur. Importantly, this distinction persists across the short-range to long-range crossover, indicating that the crossing-count diagnostic captures robust edge-versus-bulk physics rather than being tied to fine-tuned model features.